Circular No. 7525

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/1975 J 1 = S/2000 J 1 S. S. Sheppard, D. C. Jewitt, Y. Fernandez and G. Magnier, University

of Hawaii, have reported observations with the 2.2-m reflector during Nov. 21–25 of what appeared to be a new satellite of Jupiter, to be designated S/2000 J 1. While attempting to compute an orbit, B. G. Marsden, Center for Astrophysics (CfA), found a possible linkage to some observations made by M. Holman, CfA (as part of the collaboration listed on IAUC 7512), with the 3.6-m Canada-France-Hawaii telescope on Aug. 6 of what had also been considered a possible satellite. The resulting orbit was direct, but substantially smaller and more inclined than those of the other direct, outer satellites. Selected observations follow:

2000 UT α_{2000} δ_{2000} R Observer
Aug. 6.56371 4^{h}20^{m}23.^{s}30 +20^{o}57′04′′.1 20.5 Holman
Nov. 21.50367 4 23 35.29 +20 41 36.5 21.5 Sheppard
22.28860 4 23 05.01 +20 41 32.7 21.5  _{”}
23.26817 4 22 26.55 +20 41 26.9 21  _{”}
25.46439 4 20 59.01 +20 41 09.0  _{”}

Marsden also suggested the possibility of identification with the lost jovian satellite S/1975 J 1 observed by C. T. Kowal and E. Roemer (cf. IAUC 2845, 2855, 2899), this being proven by G. V. Williams, CfA, from the subarcmin agreement of a backward integration of his perturbed 2000 orbit solution. The following perturbed orbital elements by Williams satisfy 23 observations 1975 Sept. 30–2000 Nov. 25 with a mean residual of 0′′.4:

Epoch = 2000 Sept. 13.0 TT

T = 2000 Oct. 13.8356 TT  ω = 238.^{o}1484
e = 0.204340  Ω = 202.3927 }2000.0
q = 0.039290 AU  i = 45.6723

a = 0.049380 AU n^{o} = 2.7754070 P = 129.71 days

2000/01 α_{2000}  δ_{2000}  Δ  r_{J}  ϵ Δα′′  Δδ′′
Nov. 22 4^{h}23.^{m}27 +20^{o}41.^{′}5 4.011 0.055 172.^{o}5 1736+ 481+
Dec. 2 4 16.56 +20 39.4 4.004 0.058 175.8 917+ 1116+
12 4 09.86 +20 34.2 4.036 0.059 164.1  61− 1557+
22 4 03.82 +20 26.2 4.106 0.059 152.5 996− 1750+
Jan. 1 3 59.00 +20 16.3 4.210 0.057 141.2 1724− 1682+
11 3 55.80 +20 05.7 4.341 0.054 130.2 2120− 1367+
21 3 54.46 +19 55.5 4.494 0.050 119.7 2098− 841+
31 3 55.09 +19 47.1 4.658 0.045 109.6 1617− 177+
2000 November 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden